OpenStudy (anonymous):

If D varies directly with t and D is 108 when t is 4, find D when t is 7.

OpenStudy (lgbasallote):

D varies directly with t so D = kt D is 108 when t is 4 so 108 = 4k k = 108/4 k = 27 find D when t is 7 D = kt D = (28)(7) can you solve this now?

OpenStudy (lgbasallote):

did you get that @katiebug0105 ?

OpenStudy (anonymous):

Yes! Thank you!

OpenStudy (lgbasallote):

you're welcome! :)